# QGFN: Controllable Greediness with Action Values

### v1

@@ Abstract
Generative Flow Networks (GFlowNets; GFNs) are a family of reward/energy-based generative methods for combinatorial objects, capable of generating diverse and high-utility samples. However, biasing GFNs towards producing high-utility samples is non-trivial. In this work, we leverage connections between GFNs and reinforcement learning (RL) and propose to combine the GFN policy with an action-value estimate, $Q$, to create greedier sampling policies which can be controlled by a mixing parameter. We show that several variants of the proposed method, QGFN, are able to improve on the number of high-reward samples generated in a variety of tasks without sacrificing diversity.
Machine Learning, ICML
@@ 1 Introduction
Generative Flow Networks (Bengio et al., 2021a, b), also known as GFlowNets, or GFNs, were recently introduced as a novel generative framework in the family of energy based models (Malkin et al., 2022b; Zhang et al., 2022). Given some energy function $f(x)$ over objects $\mathcal{X}$, the promise of GFNs is to train a sampler $p_{\theta}$ such that at convergence $p_{\theta}(x)\propto\exp(-f(x))$; $\exp(-f(x))$ is also referred to as the reward $R(x)$ in GFN literature, inheriting terminology from Reinforcement Learning (RL). GFNs achieve this sampling via a constructive approach, treating the creation of some object $x$ as a sequential additive process (rather than an iterative local process à la Markov chain Monte Carlo (MCMC)). The main advantage of a GFN is its ability to generate a greater diversity of low-energy/high-reward objects compared to approaches based on MCMC or RL (Bengio et al., 2021a; Jain et al., 2022, 2023), or even Soft-RL (which, while related to GFNs, accomplishes something different by default, as discussed in Tiapkin et al., 2023).
@@
An important control knob in GFNs has been the use of a temperature (hyper)parameter. To get more interesting samples, it is common to train a model to sample in proportion to $R(x)^{\beta}$; $\beta$ is an inverse temperature, typically $\gg 1$, which pushes the model to become greedier. Tuning this hyperparameter is non-trivial, making it challenging to train some GFNs; for example, trajectory balance (Malkin et al., 2022a) is sensitive to the “peakiness” of the reward landscape (Madan et al., 2023). Although it is possible to train temperature-conditional models (Kim et al., 2023), doing so requires a neural network to essentially learn a whole family of GFNs–no easy task, albeit doable, e.g., in multiobjective settings (Jain et al., 2023).
@@
In this work, we propose a simple yet effective approach that allows selecting arbitrary greediness at inference time, which preserves the generality of temperature-conditionals, but can be trained in an easier way. We do so without the cost of learning a complicated family of functions and without conditionals, instead only training two models: a GFlowNet and an action-value function $Q$ (Watkins & Dayan, 1992; Mnih et al., 2013).
@@
Armed with the forward policy of a GFN, $P_{F}$ (which decides a probability distribution over actions given the current state), and the action-value, $Q$, we show that it is possible to create a variety of controllably greedy sampling policies, which can be controlled by parameters that require no retraining. We show that it is possible to simultaneously learn $P_{F}$ and $Q$, and in doing so, to generate more high-reward yet diverse object sets. In particular, we introduce and benchmark three specific variants of our approach, which we call QGFN: $p$-greedy, $p$-quantile, and $p$-of-max.
@@
We evaluate the proposed methods on 5 standard tasks used in prior GFN works: the fragment-based molecular design task introduced by Bengio et al. (2021a), 2 RNA design tasks introduced by Sinai et al. (2020), a small molecule design task based on QM9 (Jain et al., 2023), as well as a bit sequence task from Malkin et al. (2022a); Shen et al. (2023). The proposed method achieves higher average rewards and discovers modes more efficiently, sometimes by a large margin, than strong baselines.
We conduct an analysis of the proposed methods, investigate key design choices, and probe the methods to understand why they work. We also investigate other possible combinations of $Q$ and $P_{F}$–again, entirely possible at inference time, by combining trained $Q$ and $P_{F}$ models.
2 Background
@@
We follow the general setting of previous GFN literature and consider the generation of discrete finite objects, but note that in principle our method could be extended to the continuous case (Lahlou et al., 2023).
GFlowNets
@@
For a thorough treatment of GFNs, see Bengio et al. (2021b). GFNs sample objects by decomposing their generation process in a sequence $\tau=(s_{0},..,s_{T})$ of constructive steps. The space can be described by a pointed directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $s\in\mathcal{S}$ is a partially constructed object, and $(s\to s^{\prime})\in\mathcal{A}\subset\mathcal{S}\times\mathcal{S}$ is a valid additive step (e.g., adding a fragment to a molecule). $\mathcal{G}$ is rooted at a unique initial state $s_{0}$.
@@
where $P_{F}$ and $P_{B}$, the forward and backward policies, are distributions over children and parents respectively, representing the fraction of flow emanating forward and backward from a state. By construction for terminal (leaf) states $F(s)=R(s)$. Another way to think of the above is via edge flows (Bengio et al., 2021a), $F(s\to s^{\prime})=F(s)P_{F}(s^{\prime}|s)$. Again, flow conditions preserve incoming flows and outgoing flows for all states $s\in\mathcal{S}$:
$\sum_{s_{i}\in\mathrm{Parents}(s)}F(s_{i}\to s)=\sum_{s_{o}\in\mathrm{Children}(s)}F(s\to s_{o})$ (2)
where, by construction, the edge flow $F(s\to s_{T})$ to a terminal (sink) state $s_{T}$ is $R(s_{T})$ (e.g. the flow corresponding to taking the stop action), and the initial state $s_{0}$ (which has no parents) only has to account for the flow of its children (it is a source in the network).
Flow conditions lead to learning objectives such as Trajectory Balance (TB; Malkin et al., 2022a), where $n=0$ and $m$ is the trajectory length, and Sub-trajectory Balance (SubTB; Madan et al., 2023), where all combinations of $(n,m)$ are used (in a way reminiscent of TD($\lambda$); Sutton, 1988). While a variety of GFN objectives exist, we use these two standard baselines. If flow conditions are fully satisfied, i.e. 0-loss everywhere, terminal states are guaranteed to be sampled with probability $\propto R(s)$ (Bengio et al., 2021a).
Note that, by construction, the relationship between $F$ and $P_{F}$ is such that, if $s\to s^{\prime}\in\mathcal{A}$, then $P_{F}(s^{\prime}|s)=P_{B}(s|s^{\prime})F(s^{\prime})/F(s)\propto F(s^{\prime})$; the likelihood of going from $s$ to $s^{\prime}$ is proportional to the flow in $s^{\prime}$.
Action values
For a broad overview of RL, see Sutton & Barto (2018). A central object in RL is the action-value function $Q^{\pi}(s,a)$, which estimates the expected “reward-to-go” when following a policy $\pi$ starting in some state $s$ and taking action $a$; for some discount $0\leq\gamma\leq 1$,
@@
Note that $T(s,a)$ is a stochastic transition operator, but that in a GFN context objects are typically constructed in a deterministic setting (although there are stochastic extensions; Pan et al., 2023). Because rewards are only available for finished objects, $R(s)=0$ unless $s$ is a terminal state, and we use $\gamma=1$ to avoid arbitrarily penalizing “larger” objects. Finally, as there are several possible choices for $\pi$, we will simply refer to $Q^{\pi}$ as $Q$ when statements apply to a large number of such choices.
@@ 3 Motivation
Consider the following scenario, illustrated in Figure 1: an agent is faced with two doors. Behind the left door, there are 100 other doors, each hiding a reward of 1. Behind the right door, there is a single door hiding a reward of 100. The flow will be such that $F(\mathrm{left})=F(\mathrm{right})=100$, meaning that a GFN agent will pick either door with probability $\nicefrac{{1}}{{2}}$. On the other hand, the action value function is $Q(s_{0},\mathrm{left})=1$, $Q(s_{0},\mathrm{right})=100$, so an agent basing its decisions on $Q$ will reach the door with reward 100.
Figure 1: Illustration showing why solely relying on flow functions $F$ in GFNs can be insufficient. While GFNs capture how much stuff there is, they spend time sampling from lots of small rewards.
The above example shows that relying solely on flows is not always sufficient to provide high-value samples frequently and is illustrative of real-world scenarios. Consider molecular design (a very large search space) with some goodness reward in $[0,1]$; there may be $10^{6}$ molecules with reward .9.9.9.9, but just a dozen with reward $1$. Since .$9\times 10^{6}.$ is much bigger than $12\times 1$, the probability of sampling a reward 1 molecule will be low if one uses this reward naively. While using a temperature parameter is a useful way to increase the probability of the reward $1$ molecules, we propose a complementary, inference-time-adjustable method.
However, relying solely on $Q$ is also insufficient. If $Q$ were estimated very well for the optimal policy (which is extremely hard), it would be (somewhat) easy to find the reward 1 molecules via some kind of tree search following $Q$ values. However, in practice, RL algorithms easily collapse to non-diverse solutions, only discovering a few high reward outcomes. This is where flows are useful: because they capture how much stuff there is in a particular branch (rather than an expectation), it is useful to follow flows to find regions where there is potential for reward. In this paper we propose a method that can be greedier (by following $Q$) while still being exploratory and diverse (by following $F$ through $P_{F}$).
A commonly used trick in GFNs to increase the average reward, aka “greediness” of the generated samples, is to adapt the reward distribution by using an altered reward function $\hat{R}(x)=R(x)^{\beta}$ and adjusting the exponent parameter $\beta$: the higher the $\beta$, the greedier the model should be (Jain et al., 2023). However, increasing $\beta$ induces greater numerical instability. Even on a log scale, using e.g. $\beta=10$ scales a mean squared loss more or less 100×100$\times$100 ×. Another common pitfall is that increasing $\beta$ reduces the diversity of the model, because the model is less incentivized to explore “middle-reward” regions. This can lead to mode collapse. Kim et al. (2023) show that it is possible to train models that are conditioned on $\beta$, which somewhat alleviates these issues, but at the cost of training a more complex model, possibly restricted to values of $\beta$ seen during training.
Another possible approach is to leverage the equivalence between the GFN framework and the Soft-RL framework (Tiapkin et al., 2023). However, this approach produces a soft policy, while we are interested in increasing the greediness of the policy.
@@ 4 QGFN: controllable greediness through $Q$
Leveraging the intuition from the example above, we now propose and investigate several ways in which we can use $Q$-functions to achieve our goal; we call this general idea QGFN. In particular, we present three variants of this idea, which are easy to implement and effective: $p$-greedy QGFNs, $p$-quantile QGFNs, and $p$-of-max QGFNs. In §5 and §6, we show that these approaches provide a favourable trade-off between reward are diversity of the generated samples, during both training and inference.
As is common in GFlowNets, we train QGFN by sampling data from some behavior policy $\mu$. We train $F$ and $P_{F}$ (and use a uniform $P_{B}$) to minimize a flow consistency loss on the minibatch of sampled data. Additionally, we train a $Q$-network to predict action values on the same minibatch (the choice of loss will be detailed later). Training the GFN and $Q$ on a variety of behaviors $\mu$ is possible because both are off-policy methods. Indeed, instead of choosing $\mu$ to be a noisy version of $P_{F}$ as is usual for GFNs, we combine the predictions of $P_{F}$ and $Q$ to form a greedier behavior policy. In all proposed variants, this combination is modulated by a factor $p\in[0,1]$, where $p=0$ means that $\mu$ depends only on $P_{F}$, and $p=1$ means $\mu$ is greediest, as reflected by $Q$. The variants differ in the details of this combination.
$p$-greedy QGFN
Here, we define $\mu$ as a mixture between $P_{F}$ and the $Q$-greedy policy, controlled by factor $p$:
@@
$p$-quantile QGFN
Here, we define $\mu$ as a masked version of $P_{F}$, where actions below the $p$-quantile of $Q$, denoted $q_{p}(Q,s)$, have probability 00 (so are discarded):
@@
$p$-of-max QGFN
Here, we define $\mu$ as a masked version of $P_{F}$, where actions with $Q$-values less than $p\max_{a}Q(s,a)$ have probability 0:
@@
This is similar to $p$-quantile pruning, but the number of pruned actions changes as a function of $Q$. If all actions are estimated as good enough, it may be that no action is pruned, and vice versa, only the best action may be retained is none of the others are good.. This method also prunes the search space, and $p$ remains changeable. Note that in a GFN context, rewards are strictly positive, so $Q$ is also positive.
Policy evaluation, or optimal control?
@@
In the design of the proposed method, we were faced with an interesting choice: what policy $\pi$ should $Q^{\pi}$ evaluate? The first obvious choice is to perform $Q$-learning (Mnih et al., 2013), and estimate the optimal value function $Q^{*}$, with a 1-step TD objective. As we detail in §6, this proved to be fairly hard, and in general, the learned $Q_{\theta}$ ended up being a poor approximation.
@@
A commonly used trick to improve the performance of bootstrapping algorithms is to use $n$-step returns (Hessel et al., 2018). This proved essential to our work, and also revealed something curious: we’ve consistently found that, while results started improving at $n\geq 5$, a consistently good value of $n$ was close to the maximum trajectory length. This has an interesting interpretation, as beyond a certain value of $n$, $Q$ becomes closer to $Q^{\mu}$ and further from $Q^{*}$. In other words, using an “on-policy” estimate $Q^{\mu}$ rather than an estimate of the optimal policy seems beneficial, or at least easier to learn as a whole. In hindsight, this makes sense because on-policy evaluation is easier than learning $Q^{*}$, and since we are combining the $Q$-values with $P_{F}$, any method which biases $\mu$ correctly towards better actions is sufficient (we do not need to know exactly the best action, or its exact value).
Selecting greediness
@@
In the methods proposed above, $p$ can be changed arbitrarily. We first note that we train with a constant or annealed [^1] Specifically for $p$-quantile QGFN and $p$-of-max, we found training to be more stable if we started with $p=0$ and annealed $p$ towards its final value with a single $\nicefrac{{1}}{{2}}$-period cosine schedule over 1500 steps. value of $p$ and treat it as a hyperparameter. Second, as discussed in §6, after training, $p$ can be changed with a predictable effect: the closer $p$ is to 1, the greedier $\mu$ becomes. Presumably, this happens because the model is generalizing, and $Q$-value estimates for “off-policy” actions are still a reliable guess of the reward obtainable down some particular branch. When making $p$ higher, $Q$ may remain a good lower bound of the expected reward (after all, $\mu$ is becoming greedier), which is still helpful. Generally, such a policy will have reasonable outcomes, regardless of the specific $\mu$ and $p$ used during training Finally, we note it may be possible, and desirable, to use more complex schedules for $p$, or to sample $p$ during training from some (adaptive) distribution.

### v3

@@ Abstract
Generative Flow Networks (GFlowNets; GFNs) are a family of energy-based generative methods for combinatorial objects, capable of generating diverse and high-utility samples. However, consistently biasing GFNs towards producing high-utility samples is non-trivial. In this work, we leverage connections between GFNs and reinforcement learning (RL) and propose to combine the GFN policy with an action-value estimate, $Q$, to create greedier sampling policies which can be controlled by a mixing parameter. We show that several variants of the proposed method, QGFN, are able to improve on the number of high-reward samples generated in a variety of tasks without sacrificing diversity.
@@ 1 Introduction
Generative Flow Networks (Bengio et al., 2021a, b), also known as GFlowNets, or GFNs, were recently introduced as a novel generative framework in the family of energy-based models (Malkin et al., 2022b; Zhang et al., 2022). Given some energy function $f(x)$ over objects $\mathcal{X}$, the promise of GFNs is to train a sampler $p_{\theta}$ such that at convergence $p_{\theta}(x)\propto\exp(-f(x))$; $\exp(-f(x))$ is also referred to as the reward $R(x)$ in GFN literature, inheriting terminology from Reinforcement Learning (RL). GFNs achieve this sampling via a constructive approach, treating the creation of some object $x$ as a sequential additive process (rather than an iterative local process à la Markov chain Monte Carlo (MCMC) that can suffer from mode-mixing issues). The main advantage of a GFN is its ability to generate a greater diversity of low-energy/high-reward objects compared to approaches based on MCMC or RL (Bengio et al., 2021a; Jain et al., 2022, 2023), or even Soft-RL–which, while related to GFNs, accomplishes something different by default (Tiapkin et al., 2023; Mohammadpour et al., 2024; Deleu et al., 2024).
@@
To generate more interesting samples and avoid oversampling from low-reward regions, it is common to train a model to sample in proportion to $R(x)^{\beta}$; $\beta$ is an inverse temperature, typically $\gg 1$, which pushes the model to become greedier. The use of this temperature (hyper)parameter is an important control knob in GFNs. However, tuning this hyperparameter is non-trivial, which complicates training certain GFNs; for example, trajectory balance (Malkin et al., 2022a) is sensitive to the “peakiness" of the reward landscape (Madan et al., 2023). Although it is possible to train temperature-conditional models (Kim et al., 2023a), doing so essentially requires learning a whole family of GFNs–no easy task, albeit doable, e.g., in multiobjective settings (Jain et al., 2023).
@@
In this work, we propose an approach that allows selecting arbitrary greediness at inference time, which preserves the generality of temperature-conditionals, while simplifying the training process. We do so without the cost of learning a complicated family of functions and without conditionals, instead only training two models: a GFlowNet and an action-value function $Q$ (Watkins and Dayan, 1992; Mnih et al., 2013).
@@
Armed with the forward policy of a GFN, $P_{F}$ (which decides a probability distribution over actions given the current state, i.e. $\pi$ in RL), and the action-value, $Q$, we show that it is possible to create a variety of controllably greedy sampling policies, controlled by parameters that require no retraining. We show that it is possible to simultaneously learn $P_{F}$ and $Q$, and in doing so, to generate more high-reward yet diverse object sets. In particular, we introduce and benchmark three specific variants of our approach, which we call QGFN: $p$-greedy, $p$-quantile, and $p$-of-max.
@@
We evaluate the proposed methods on 5 standard tasks used in prior GFN works: the fragment-based molecular design task introduced by Bengio et al. (2021a), 2 RNA design tasks introduced by Sinai et al. (2020), a small molecule design task based on QM9 (Jain et al., 2023), as well as a bit sequence task from Malkin et al. (2022a); Shen et al. (2023). The proposed method outperforms strong baselines, achieving high average rewards and discovering modes more efficiently, sometimes by a large margin. We conduct an analysis of the proposed methods, investigate key design choices, and probe the methods to understand why they work. We also investigate other possible combinations of $Q$ and $P_{F}$–again, entirely possible at inference time, by combining trained $Q$ and $P_{F}$ models.
2 Background and Related Work
@@
We follow the general setting of previous GFN literature and consider the generation of discrete finite objects, but in principle our method could be extended to the continuous case (Lahlou et al., 2023).
@@
GFlowNets    GFNs (Bengio et al., 2021b) sample objects by decomposing their generation process in a sequence $\tau=(s_{0},..,s_{T})$ of constructive steps. The space can be described by a pointed directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $s\in\mathcal{S}$ is a partially constructed object, and $(s\to s^{\prime})\in\mathcal{A}\subset\mathcal{S}\times\mathcal{S}$ is a valid additive step (e.g., adding a fragment to a molecule). $\mathcal{G}$ is rooted at a unique initial state $s_{0}$.
@@
where $P_{F}$ and $P_{B}$, the forward and backward policies, are distributions over children and parents respectively, representing the fraction of flow emanating forward and backward from a state. By construction for terminal (leaf) states $F(s)=R(s)$.
Balance conditions lead to learning objectives such as Trajectory Balance (TB; Malkin et al., 2022a), where $n=0$ and $m$ is the trajectory length, and Sub-trajectory Balance (SubTB; Madan et al., 2023), where all combinations of $(n,m)$ are used. While a variety of GFN objectives exist, we use these two as they are considered standard. If those objectives are fully satisfied, i.e. 0-loss everywhere, terminal states are guaranteed to be sampled with probability $\propto R(s)$ (Bengio et al., 2021a).
Action values    For a broad overview of RL, see Sutton and Barto (2018). A central object in RL is the action-value function $Q^{\pi}(s,a)$, which estimates the expected “reward-to-go” when following a policy $\pi$ starting in some state $s$ and taking action $a$; for some discount $0\leq\gamma\leq 1$,
@@
While $T(s,a)$ can be a stochastic transition operator, in a GFN context objects are constructed in a deterministic way (although there are stochastic GFN extensions; Pan et al. (2023b)).
Because rewards are only available for finished objects, $R(s)=0$ unless $s$ is a terminal state, and we use $\gamma=1$ to avoid arbitrarily penalizing “larger” objects. Finally, as there are several possible choices for $\pi$, we will simply refer to $Q^{\pi}$ as $Q$ when statements apply to a large number of such choices.
2.1 Related Work
RL and GFlowNets    There are clear connections between the GFN framework and RL framework (Tiapkin et al., 2023; Mohammadpour et al., 2024; Deleu et al., 2024). Notably, Tiapkin et al. (2023) show that it is possible to reformulate fixed-$P_{B}$ GFlowNets as a Soft-RL problem within a specific class of reward-modified MDPs. While they show that this reformulated problem can then be tackled with any Soft-RL method, this still essentially solves the original GFlowNet problem, i.e. learn $p_{\theta}(x)\propto R(x)$. Instead, we are interested in greedier-yet-diverse methods.
A commonly used tool in GFNs (and QGFN) to increase the average reward, aka “greediness" of the drawn samples, is to adapt the reward distribution by using an altered reward function $\hat{R}(x)=R(x)^{\beta}$ and adjusting the exponent parameter $\beta$: the higher the $\beta$, the greedier the model should be (Jain et al., 2023). However, increasing $\beta$ often induces greater numerical instability (even on a log scale), and reduces diversity because the model is less incentivized to explore “middle-reward” regions. This can lead to mode collapse. Kim et al. (2023a) show that it is possible to train models that are conditioned on $\beta$, which somewhat alleviates these issues, but at the cost of training a more complex model.
Again, while we could leverage the equivalence between the GFN framework and the Soft-RL framework (Tiapkin et al., 2023), this approach would produce a soft policy. We propose a different approach that increases greediness of the policy via “Hard-RL”.
Improving GFlowNet sampling    A number of works have also made contributions towards improving utility in GFlowNets, via local search (Kim et al., 2023b), utilizing intermediate signals (Pan et al., 2023a), or favoring high-potential-utility intermediate states (Shen et al., 2023), as well as the use of RL tools such as replay buffers (Vemgal et al., 2023), target networks (Lau et al., 2023), or Thompson sampling (Rector-Brooks et al., 2023).
@@ 3 Motivation
Figure 1: Solely relying on flow functions $F$ in GFNs can be insufficient. While GFNs capture how much stuff there is, they spend time sampling from lots of small rewards.
Consider the following scenario, illustrated in Figure 1: an agent is faced with two doors. Behind the left door, there are 100 other doors, each hiding a reward of 1. Behind the right door, there is a single door hiding a reward of 100. The flow will be such that $F(\mathrm{left})=F(\mathrm{right})=100$, meaning that a GFN agent will pick either door with probability $\nicefrac{{1}}{{2}}$. The action value function is $Q(s_{0},\mathrm{left})=1$, $Q(s_{0},\mathrm{right})=100$, so an agent basing its decisions on $Q$ will reach for the door with reward 100.
This example shows that relying solely on flows is not always sufficient to provide high-value samples frequently and is illustrative of real-world scenarios. Consider molecular design (a very large search space) with some reward in $[0,1]$; there may be $10^{6}$ molecules with reward .9.9.9.9, but just a dozen with reward $1$. Since .$9\times 10^{6}.$ is much bigger than $12\times 1$, the probability of sampling a reward 1 molecule will be low if one uses this reward naively. While using a temperature parameter is a useful way to increase the probability of the reward $1$ molecules, we propose a complementary, inference-time-adjustable method.
Note that relying solely on $Q$ is also insufficient. If $Q$ were estimated very well for the optimal policy (which is extremely hard), it would be (somewhat) easy to find the reward 1 molecules via some kind of tree search following $Q$ values. However, in practice, RL algorithms easily collapse to non-diverse solutions, only discovering a few high reward outcomes. This is where flows are useful: because they capture how much stuff there is in a particular branch (rather than an expectation), it is useful to follow flows to find regions where there is potential for reward. In this paper, we propose a method that can be greedier (by following $Q$) while still being exploratory and diverse (by following $F$ through $P_{F}$).
@@ 4 QGFN: controllable greediness through $Q$
Leveraging the intuition from the example above, we now propose and investigate several ways in which we can use $Q$-functions to achieve our goal; we call this general idea QGFN. In particular, we present three variants of this idea, which are easy to implement and effective: $p$-greedy QGFNs, $p$-quantile QGFNs, and $p$-of-max QGFNs. In §5 and §6, we show that these approaches provide a favourable trade-off between reward and diversity, during both training and inference.
As is common in GFlowNets, we train QGFN by sampling data from some behavior policy $\mu$. We train $F$ and $P_{F}$ (and use a uniform $P_{B}$) to minimize a flow balance loss on the minibatch of sampled data, using a temperature parameter $\beta$. Additionally, we train a $Q$-network to predict action values on the same minibatch (the choice of loss will be detailed later). Training the GFN and $Q$ on a variety of behaviors $\mu$ is possible because both are off-policy methods. Indeed, instead of choosing $\mu$ to be a noisy version of $P_{F}$ as is usual for GFNs, we combine the predictions of $P_{F}$ and $Q$ to form a greedier behavior policy. In all proposed variants, this combination is modulated by a factor $p\in[0,1]$, where $p=0$ means that $\mu$ depends only on $P_{F}$, and $p=1$ means $\mu$ is greediest, as reflected by $Q$. The variants differ in the details of this combination.
$p$-greedy QGFN    Here, we define $\mu$ as a mixture between $P_{F}$ and the $Q$-greedy policy, controlled by factor $p$:
@@
$p$-quantile QGFN    Here, we define $\mu$ as a masked version of $P_{F}$, where actions below the $p$-quantile of $Q$, denoted $q_{p}(Q,s)$, have probability 00 (so are discarded):
@@
$p$-of-max QGFN    Here, we define $\mu$ as a masked version of $P_{F}$, where actions with $Q$-values less than $p\max_{a}Q(s,a)$ have probability 0:
@@
This is similar to $p$-quantile pruning, but the number of pruned actions changes as a function of $Q$. If all actions are estimated as good enough, it may be that no action is pruned, and vice versa, only the best action may be retained is none of the others are good.
This method also prunes the search space, and $p$ remains changeable. Note that in a GFN context, rewards are strictly positive, so $Q$ is also positive.
@@
Policy evaluation, or optimal control?    In the design of the proposed method, we were faced with an interesting choice: what policy $\pi$ should $Q^{\pi}$ evaluate? The first obvious choice is to perform $Q$-learning (Mnih et al., 2013), and estimate the optimal value function $Q^{*}$, with a 1-step TD objective. As we detail in §6, this proved to be fairly hard, and 1-step $Q_{\theta}$ ended up being a poor approximation.
@@
A commonly used trick to improve the performance of bootstrapping algorithms is to use $n$-step returns (Hessel et al., 2018). This proved essential to our work, and also revealed something curious: we’ve consistently found that, while results started improving at $n\geq 5$, a consistently good value of $n$ was close to the maximum trajectory length. This has an interesting interpretation, as beyond a certain value of $n$, $Q_{\theta}$ becomes closer to $Q^{\mu}$ and further from $Q^{*}$. In other words, using an “on-policy” estimate $Q^{\mu}$ rather than an estimate of the optimal policy seems beneficial, or at least easier to learn as a whole. In hindsight, this makes sense because on-policy evaluation is easier than learning $Q^{*}$, and since we are combining the $Q$-values with $P_{F}$, any method which biases $\mu$ correctly towards better actions is sufficient (we do not need to know exactly the best action, or its exact value).
Selecting greediness    In the methods proposed above, $p$ can be changed arbitrarily. We first note that we train with a constant or annealed [^2] Specifically for $p$-quantile QGFN and $p$-of-max, we found training to be more stable if we started with $p=0$ and annealed $p$ towards its final value with a single $\nicefrac{{1}}{{2}}$-period cosine schedule over 1500 steps. value of $p$ and treat it as a standard hyperparameter in all the results reported in §5.
@@
Second, as discussed in §6, after training, $p$ can be changed with a predictable effect: the closer $p$ is to 1, the greedier $\mu$ becomes. Presumably, this is because the model generalizes, and $Q$-value estimates for “off-policy” actions are still a reliable guess of the reward obtainable down some particular branch. When making $p$ higher, $Q$ may remain a good lower bound of the expected reward (after all, $\mu$ is becoming greedier), which is still helpful.
Generally, such a policy will have reasonable outcomes, regardless of the specific $\mu$ and $p$ used during training. Finally, it may be possible and desirable to use more complex schedules for $p$, or to sample $p$ during training from some (adaptive) distribution, but we leave this for future work.
Figure 2: Fragment-based molecule task. Left: Average rewards over the training trajectories. Center: Number of unique modes with a reward threshold exceeding 0.97 and pairwise Tanimoto similarity score less than 0.65. Right: Average pairwise Tanimoto similarity score for the top 1000 molecules sampled by reward. Lines are the interquartile mean and standard error calculated over 5 seeds.